OpenStudy (anonymous):

Use natural logarithms to solve the equation. Round to the nearest thousand. 3e^4x-12=29. HAve no idea how to solve this, help?

OpenStudy (campbell_st):

is the equation. 1. \[3e^{4x -12} = 29\] or 2. \[3e^{4x} - 12 = 29\]

OpenStudy (anonymous):

@campbell_st number 2 :)

OpenStudy (campbell_st):

ok... start by adding 12 to both sides of the equation \[3e^{4x} = 41\] next divide by 3 \[e^{4x} = \frac{41}{3}\] now take the ln of both sides \[4x = \ln(\frac{41}{3})\] now solve for x

OpenStudy (anonymous):

@campbell_st so you divide both numbers by 4 and they come out as a decimal correct?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

@campbell_st idk if i divided it correctly but would it be 3.4166?

OpenStudy (campbell_st):

well you seem to have an error... what is ln(41/3) = then divide that answer by 4

OpenStudy (anonymous):

@campbell_st so ln(43/3) is 2.61495? so then it would equal 0.6537?

OpenStudy (campbell_st):

yep... thats what I calculated x to be... you can check it by substitution \[3e^{4 \times 0.6537} - 12 = 29\]

OpenStudy (anonymous):

@campbell_st thanks so much you helped me out a lot!

OpenStudy (campbell_st):

no problem... just make sure you don't use logs to early in this type of question... good luck